A 780-mile, 7-hour plane trip was flown at two speeds. For the first part of the trip, the average speed was 105 mph.
For the remainder of the trip, the average speed was 120 mph. How long did the plane fly at each speed?
105 mph

120 mph

Answer by ikleyn(52790)

.
A 780-mile, 7-hour plane trip was flown at two speeds. For the first part of the trip, the average speed was 105 mph.
For the remainder of the trip, the average speed was 120 mph. How long did the plane fly at each speed?
105 mph

120 mph
~~~~~~~~~~~~~~~~~~~~~~~~~

Solution 1 (system of two equations)

   x +     y =  7,       (1)    (x is time flying at 105 mph; y is time flying at 120 mph)
105x + 120y = 780.       (2)

Express x = 7-y from (1) and substitute into (2), then solve or y.

Solution 2 (one equation)

105*x _ 120*(7-x) = 780.

Solve for x.